OpenStudy (anonymous):

Prove the following identity: sec(2θ) = sec^2 (θ)/2-sec^2(θ)

OpenStudy (anonymous):

\(\sec\theta=1/\cos\theta\) then use the sum identity for cos

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

you're welcome!